OpenStudy (anonymous):

ILL GIVE YOU MEDALS Find the least common multiple of : 16x3yz3 and 14z6xy4c

OpenStudy (anonymous):

Lol finishing in two minutes. Really XD What is the lowest number that can go in both 16 and 14?

OpenStudy (anonymous):

would the answer be 2xyz^3 ?

OpenStudy (anonymous):

@KirbyLegs

OpenStudy (ranga):

Do it term by corresponding term: What is LCM of 16 and 14? LCM of x^3 and x LCM of y and y^4 LCM of Z^3 (Is the last term c or z?)

OpenStudy (anonymous):

the last term is z

OpenStudy (whpalmer4):

least common multiple is found by factoring each of the items, then multiplying the highest power found of each unique factor. LCM of 8 and 9, for example: 8 = 2*2*2 = 2^3 9 = 3*3 = 3^2 2^3*3^2 = 8*9 = 72 check: 8 16 24 32 40 48 56 64 72 80 88 96 9 18 27 36 45 54 63 72 81 90 72 is indeed the first common multiple in the two lists

OpenStudy (anonymous):

Medal fishing can get you suspended